OpenStudy (anonymous):

3x-4y=1 in simpilest form

OpenStudy (anonymous):

Simplifying 3x + -4y = 1 Solving 3x + -4y = 1 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '4y' to each side of the equation. 3x + -4y + 4y = 1 + 4y Combine like terms: -4y + 4y = 0 3x + 0 = 1 + 4y 3x = 1 + 4y Divide each side by '3'. x = 0.3333333333 + 1.333333333y Simplifying x = 0.3333333333 + 1.333333333y Simplifying 3x + -4y = 1 Solving 3x + -4y = 1 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '4y' to each side of the equation. 3x + -4y + 4y = 1 + 4y Combine like terms: -4y + 4y = 0 3x + 0 = 1 + 4y 3x = 1 + 4y Divide each side by '3'. x = 0.3333333333 + 1.333333333y Simplifying x = 0.3333333333 + 1.333333333y Simplifying x = 0.3333333333 + 1.333333333y